OpenStudy (anonymous):

Find the sum of the first 125 positive even whole numbers.

OpenStudy (anonymous):

if 0 is excluded; then we have 2,4,6,8..forming an AP. then 1st term A = 2 , common difference D = 2, and the number of terms N = 125. find the last term L = A + (N-1)D with A = 2, D = 2, N = 125. and tell me the answer.

OpenStudy (anonymous):

250

OpenStudy (anonymous):

ok. so A = 2 and L = 250. Use the formula Sum S = N X (A+L) / 2 with N = 125, A = 2, L = 250; write down d ans.

OpenStudy (anonymous):

15750